OpenStudy (summersnow8):

A 2.00 lb rock is dropped from a bridge to the water 50.0 ft below. What is its kinetic energy (in ft lb) as it hits the water. book says 99.8 lb ft @radar

OpenStudy (anonymous):

Do you know how to calculate kinetic energy?

OpenStudy (summersnow8):

@ospreytriple KE = 1/2 m v^2 ?

OpenStudy (anonymous):

That's right. The mass can be determined from the given weight of 2 lb using\[F _{g} = mg\]I assume you're OK with doing that. What's left, then is to determine the velocity of the rock when it hits the water. Knowing that acceleration is -32.2 ft/s^2 we could use\[v _{f}^{2} = v _{i}^{2} + 2a \Delta d\]where v_i = 0, a = -32.2, and d = -50

OpenStudy (anonymous):

Then, knowing m and v_f, you'll be able to calculate KE.

OpenStudy (summersnow8):

so i got Fg = 64.4 Vf = 56.7

OpenStudy (anonymous):

We agree on the Vf. But I don't use the Imperial system of units often. Does the rock have a weight of 2 lbf or a mass of 2 lbm?

OpenStudy (summersnow8):

the rock is in lbs.... the answer is supposed to be in lb ft

OpenStudy (summersnow8):

gravity for ft/s ^2 is 32.2 ft/s^2

OpenStudy (anonymous):

OK. So if the weight (Fg) of the rock is 2.00 lb, we need to rearrange the weight formula to solve for mass, i.e.\[F _{g}=mg\]\[m=\frac{ F _{g} }{ g } = \frac{ 2.00 \text{ lb} }{ 32.2 \frac{ \text{ft} }{ \text{s} ^{2}}}\]

OpenStudy (anonymous):

Now what do you get for mass?

OpenStudy (summersnow8):

0.06211 lb

OpenStudy (anonymous):

Good. If my memory is correct, the units of mass in this system are slugs. Anyway, doesn;t matter. Use this value of mass and the value of Vf you calculated previously and use the KE equation. What do you get?

OpenStudy (summersnow8):

KE = 1/2 (.06211 lb) (56.7 m/s) = 99.83

OpenStudy (anonymous):

Looks good to me. Congrats!

OpenStudy (anonymous):

One quick question before you go. I assume you're in the U.S. I'm in Canada. Do you ever work in the metric system in your physics class?

OpenStudy (summersnow8):

yes, we need to know how to convert back and forth

OpenStudy (anonymous):

OK. Thanks. And do you call these units of mass SLUGS?

OpenStudy (summersnow8):

some of the problems I have solved involved slugs, but not all of them

OpenStudy (anonymous):

Well, you've done very well. It's been great working with you.